OpenStudy (anonymous):

Binomial Expansion : (1+x)^n type...

OpenStudy (anonymous):

\[\frac{ 1 }{ 1-2x } +\frac{ 1 }{ 2+x } - \frac{ 2 }{ (2+x)^{-2} }\]

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

what class u in, and what school u go to?

OpenStudy (anonymous):

A level CIE exams for HSc.

OpenStudy (anonymous):

The answer should be : 1+ 9/4 x + 15/4 x^2 ...

OpenStudy (anonymous):

obtained : 1 + 5/2 x + -3x^2 :S

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

\[( 1 -2x)^{-1} + (2+x)^{-1} - 2(2+x)^{-2}\] is that correct ?

OpenStudy (anonymous):

\[(1+ n)^{x} = 1 + nx +\frac{ n( n-1) }{ 2! } x ^{2} +...+\]

OpenStudy (anonymous):

i need to find terms up to x^2

OpenStudy (anonymous):

After Partial Fractions im asked to find binomial expansion up to x^2 the partial fractions results is correct but the binomial expansion result is not correct someone please help :D

hero (hero):

Yeah, you should have clarified that at the beginning that way I could ping someone like @ganeshie8

OpenStudy (anonymous):

\[(1-2x)^{-1} = 1 + (-1) (-2x) + \frac{ -1(-1-1) }{ 2! }x ^{2}\]

OpenStudy (anonymous):

im not sure if this one is good how to make ( 1-2x)^-1 into ( 1+ x)^n form ?

OpenStudy (anonymous):

@hartnn can you help me with that :D?

hartnn (hartnn):

(....)^-2 <------what ? you are done with partial fractions, right ? denominators are 1-2x, x+2 and (x+2)^2 isn't it?

OpenStudy (anonymous):

done with partial fractions and the denominators are 1-2x , x+2 and (2+x)^2

hero (hero):

That's not what you posted originally @antoni7

hartnn (hartnn):

right, so \((1-2x)^{-1} = 1 + (-1) (-2x) + \frac{ -1(-1-1) }{ 2! }x ^{2}\) is absolutely correct, what further doubts u have ?

hartnn (hartnn):

wait, i think there's an error

OpenStudy (anonymous):

\[(1+\frac{ x }{ 2 }) ^{-1} = 1+ (-1) (\frac{ x }{ 2 }) + \frac{ -1(-1-1) }{ 2! }(\frac{ x }{ 2 })\]

OpenStudy (anonymous):

\[(\frac{ x }{ 2 }) ^{2}\] i

OpenStudy (anonymous):

last part is (x/2)^2 ..

hartnn (hartnn):

(1+x)^n = 1+nx+n(n-1)x^2/2 (till x^2 term) so, for (1-2x)^-1 n=-1, x=-2x 1+2x +4x^2 got this first ?

hartnn (hartnn):

it should be 4x^2 and not just x^2 you got

OpenStudy (anonymous):

yes got that

OpenStudy (anonymous):

Here is what i got for all : 1 +2x +4x^2 +1 - x/2 +x^2/4 -1 + x - 3x^2/4

hartnn (hartnn):

-2/(x+2)^2=-1/2 + x/2 -3x^2/8

hartnn (hartnn):

did u miss the 2 ? in last 3 terms ??

hartnn (hartnn):

-2/(x+2)^2= -1/2 * (1+x/2)^-2

OpenStudy (anonymous):

last term : \[- ( 1 + (-2) \frac{ x }{ 2 } + \frac{ -2 (-2-1) }{ 2! } (\frac{ x }{ 2 })^{2}\]

hartnn (hartnn):

and u forgot the initial -1/2 .....

OpenStudy (anonymous):

o.O

hartnn (hartnn):

u got this : "-2/(x+2)^2= -1/2 * (1+x/2)^-2" right ?

OpenStudy (anonymous):

\[-\frac{ 2 }{ (2+x)^{2} } = -2 (2+x)^{-2} = - (1+ \frac{ x }{ 2 })^{-2}\]

hartnn (hartnn):

no^

OpenStudy (anonymous):

is it -1/2 or -1 O.O ?

ganeshie8 (ganeshie8):

\(\large -2(2+x)^{-2} = -2 \times (2)^{-2} (\frac{2+x}{2})^{-2} = \frac{-1}{2} (1+\frac{x}{2})^{-2} \)

hartnn (hartnn):

\((x+2)^2 = (2(1+x/2))^2=2^2 (1+x/2)^2= 4 (1+x/2)^2 \)

OpenStudy (anonymous):

Omg!

hartnn (hartnn):

got it now ?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

so i got : 3/2 + 2x + 31x^2/8 and the answerr should be as above

hartnn (hartnn):

because 1st term is negative -1/(1-2x) = -1-2x-4x^2 you forgot to consider the - *minus*

OpenStudy (anonymous):

wait wait wait :D lets clear it all up :D

OpenStudy (anonymous):

\[\frac{ 4 +5x -x ^{2} }{ (1-2x) (2+x)^{2} } = \frac{ 1 }{ 1-2x } + \frac{ 1 }{ 2+x } -\frac{ 2 }{ (2+x)^{2} }\]

OpenStudy (anonymous):

\[\frac{ 4 +5x -x ^{2} }{ (1-2x) (2+x)^{2} } = \frac{ 1 }{ 1-2x } + \frac{ 1 }{ 2+x } -\frac{ 2 }{ (2+x)^{2} }\]

hartnn (hartnn):

correct

OpenStudy (anonymous):

so \[(1-2x)^{-1} + ( 1 + \frac{ x }{ 2 })^{-1} - \frac{ 1 }{ 2 } ( 1 + \frac{ x }{ 2 })^{-2}\]

OpenStudy (anonymous):

correct ?

hartnn (hartnn):

1/(x+2) =.... i think here was the error 1/(x+2) = 1/2 -x/4+x^2/8

hartnn (hartnn):

(1-2x)^-1 = 1+2x+4x^2 was correct

hartnn (hartnn):

-2/(x+2)^2=-1/2 + x/2 -3x^2/8 was correct

hartnn (hartnn):

now add

OpenStudy (anonymous):

greaatttttttt yes 100% right answer for you :D

OpenStudy (anonymous):

the problem was 1/2+x expansion.. jeez...

hartnn (hartnn):

yeah ...big problem, silly mistakes can occur, but we finally nailed it :D

OpenStudy (anonymous):

Super big thanks to all of you for correcting my silly mistakes.... :D